OpenStudy (anonymous):

what is the expected value of x^2 in mathematical form?

OpenStudy (anonymous):

in a simplified form

OpenStudy (lgbasallote):

what do you mean?

Parth (parthkohli):

\(x \times x\) I think that's what you mean.

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

mathematical form?

Parth (parthkohli):

Hmm. Similarly: \( \color{Black}{\Rightarrow x^3 = x \times x\times x}\) \( \color{Black}{\Rightarrow x^4 = x \times x \times x \times x}\)

Parth (parthkohli):

'Mathematical form' is still ambiguous.

OpenStudy (anonymous):

some stats people know may know

OpenStudy (unklerhaukus):

\[\langle x^2\rangle =\sum\limits_{j=0}^\infty x^2 P(j)\]

Parth (parthkohli):

lol

OpenStudy (lgbasallote):

\[\int 2xdx\]

OpenStudy (anonymous):

this reminds me of a very fundamental/basic question in calculus..please see my next question..i'll post link..

OpenStudy (lgbasallote):

\[\frac{d}{dx} (\frac{x^3}{3})\]

OpenStudy (unklerhaukus):

what are you doing at @lgbasallote ,

OpenStudy (lgbasallote):

\[x^2 = r^2 - y^2\]

OpenStudy (lgbasallote):

trying to get all that's x^2 and see if anythign suits him

OpenStudy (unklerhaukus):

im pretty sure i have provided the answer to the question

OpenStudy (lgbasallote):

we all think that...

OpenStudy (unklerhaukus):

,oh

OpenStudy (anonymous):

i need E(x^2)=?

OpenStudy (unklerhaukus):

the expectiation value of \(x\) \[\langle x\rangle =E(x)\]

OpenStudy (anonymous):

@UnkleRhaukus yes but x^2

OpenStudy (lalaly):

\[Var(x)=E(x^2)-(E(x))^2\]

OpenStudy (anonymous):

i got the ans... thanks all

OpenStudy (unklerhaukus):

\[\sigma_x=\langle x^2\rangle-\langle x\rangle ^2\]

OpenStudy (lalaly):

:D

OpenStudy (anonymous):

ya ur right

OpenStudy (unklerhaukus):

The expectation value , or expect value of a function is \[\langle f(x)\rangle =\sum\limits_{x=0}^\infty f(x)P(x)\] where \(P(x)\) is the probability of x

OpenStudy (anonymous):

yes for a discrete random variable

OpenStudy (unklerhaukus):

im not sure why i put j instead of x,

OpenStudy (unklerhaukus):

oh, you want a continuous function?

OpenStudy (anonymous):

no i know

OpenStudy (anonymous):

thank you

OpenStudy (unklerhaukus):

\[\langle f(x)\rangle=\int\limits_{-\infty}^\infty f(x)\rho(x)\text dx\]

OpenStudy (anonymous):

ya

OpenStudy (unklerhaukus):

where \(\rho(x)\) is the probability density

OpenStudy (unklerhaukus):

so \[\langle x^2\rangle=\int\limits_{-\infty}^\infty x^2\rho(x)\text dx\]

OpenStudy (anonymous):

@UnkleRhaukus small question the E(constant) is a constant right

OpenStudy (unklerhaukus):

if the distribution of the variable \(x\) is constant , yes

OpenStudy (anonymous):

thanks